The two forces F1 and F2 inclined at an angle of 60 degrees with each other is acting on a body of mass 5kg. If F1 = 20N and F2 = 15N, what will be the magnitude and direction of acceleration?

Grade:11

2 Answers

arnab nandy
13 years ago

x direction : f2+cos 60 xf2 =15+10=25=m A(x direction)

 

Ax in x direction is 25/5 =5m/s

 

in y 

20 sin 60 = m Ay

Ay=2(3)^(1/2)

net acc is A:

A^2 = 25+12

A=(37)^(1/2)

ROHTAZ SINGH (IIT M)
13 years ago

MAGNITUDE OF FORCE

F=( (20)2 +(15)2 ) 1/2

 F=25N

  FOR DIRECTION


Tan C    =   16sin  60/20+16cos 60

WHERE C IS ANGLE BETWEEN 20N FORCE AND RESULTANT FORCE

 

THANK YOU

ROHTAZ SINGH